# Early-Exit Graph Neural Networks

### v1

@@ E.3 Implementation Details
The metrics that we use are the Area Under the Curve (AUC) for the binary classification experiments, accuracy for Roman Empire and Amazon Ratings, average precision (AP) for Peptides-func, mean-absolute error (MAE) for Peptides-struct, and macro F1-Score (F1) for Pascal-VOC. All the experiments concerning EEGNN and SAS-GNN implied that we removed self-loops from the graphs, in order to satisfy the assumptions made in Theorems 3.1, 3.2. We always apply $f(\mathbf{X})$ implemented via a one-layer perception with ReLU. And in the LRGB, we also use a final decoder MLP to refine the representations of $\mathbf{Z}$, details on the hyperparameter are in the code. In the heterophilic graphs, each training lasts 3000 epochs, and the optimizer that we use is Adam [Kingma and Ba, 2017]. Concerning the quantitative results, we report the mean and standard deviation, which are computed from the metric averaged across 10 runs. Each run corresponds to a different split of the nodes used for training, validation, and testing. These experiments were run on a single Nvidia GeForce RTX 3090 Ti 24 GB and also a single NVIDIA A100 80GB PCIe. As concerns the LRGB experiments, we run each configuration for a maximum of 1000 epochs across 4 different weight initialization, using the AdamW optimizer. We picked the best configuration based on the validation metric. These experiments were run on a single NVIDIA A100 80GB PCIe. As concerns with the experiments on the TUDatasets for short-range graph classification (see Table 17) and homophilic node classification (see Section F.7), we also used a single NVIDIA A100 80GB PCIe. All the data was taken mostly from the PyTorch Geometric library [Fey and Lenssen, 2019], and other open-source sources. We always refer to their provenance in the code when using them.
@@ E.4 Hyperparameters
For a fair comparison, the hyperparameter tuning process was done according to the hyperparameter set used by Finkelshtein et al. [2024]. In the SAS-GNN experiments, we choose $\tau$ among $\{0.1,0.5,1\}$. While in the experiments with EEGNNs, we do not need to tune $\tau$ because it is adaptive (see Algorithm 1), and we do not need to tune the number of layers parameter since we use a large and upper bound depth $L=20$. In the case of EEGNN, we choose $f_{c}(\cdot)$ as a Mean-GNN [Hamilton, ]. Similarly to Finkelshtein et al. [2024], we test a contained number of layers $\{1,2,3\}$, and then as hidden dimension we select $4$, $8$, $16$, $32$ or $64$. Then we also tune the value of the temperature $\nu_{0}$ between 00 and 0.1. $f_{\nu}(\cdot)$, the network that learns the adaptive temperature $\nu$, adopts the design proposed in [Finkelshtein et al., 2024]. In the experiments concerning the LRGB datasets, $f_{c}$ and $f_{\nu}$ keep the same hyperparameter set as the heterophilic experiments. The main hyperparameter set is the same used in Tönshoff et al. [2023]. In any experiment with SAS-GNN and EEGNN, we do not use any dropout or normalization layer within the message-passing updates, to follow the ODE interpretation and directly minimize the energy functional in Equation (3). We only use them in the feature encoder and in the final decoder. In the experiments with Peptides-func and Peptides-struct, we use two positional encoding techniques, which are typically approached in these settings. For Peptides-func we used RWSE [Dwivedi et al., 2022], while instead in Peptides-struct, we used LapPE [Dwivedi and Bresson, 2021b]. We used these according to the best hyperparameters used for graph convolutional methods in Tönshoff et al. [2023]. We do not use positional encodings in Pascal-VOC.
@@ Appendix F Additional Results
F.1 SAS-GNN is able to retain performance as depth increases
Having proposed SAS-GNN, we have demonstrated that it is always stable and non-dissipative, and also able to allow each node to induce attraction or repulsion edge-wise. We study the behavior of SAS-GNN when the depth increases up to 50 layers, which is an overestimate of what we need in practice. In our experiments, we design SAS-GNN, as specified in Corollary 3.3. All the data points corresponds to a different training procedure, where we display the best accuracy. We present the results in Figure 4(a), where we show an experiment with Cora, a homophilic dataset. In Figure 4(b), we assess the resilience to depth via Minesweeper, an example of a heterophilic graph. In Cora, we notice that both A-DGN and SAS-GNN can retain their performance. Also, GRAFF can maintain it up to the 30th layer, then it starts decreasing. This behavior is expected since GRAFF is only designed to contrast over-smoothing, which does not occur, but it is not specifically designed to be robust to the number of layers. It is interesting to notice that SAS-GNN, even though part of its architecture is inherited by GRAFF, can retain its performance, behaving similarly to A-DGN. As far as GCN is concerned, we see that it has started losing its performance in the earlier layers. Here, we consider a GCN design that lacks the residual connection and does not use weight sharing. Furthermore, Table 8 presents an individual analysis of SAS-GNN on the Minesweeper dataset as the number of layers increases up to 100. In this experiment, we also allow the hidden dimension $m^{\prime}$ to scale to assess the robustness of the model under varying capacity. We observe that the model maintains its performance as the number of layers increases. However, increasing $m^{\prime}$ from 32 to 256 leads to a drop in performance. We attribute this decline to reduced generalization at higher hidden dimensions. Notably, $m^{\prime}=32$ emerged as the optimal setting based on our hyperparameter tuning procedure.
Table 8: Performance across different model sizes and layer depths
Model 10 layers 50 layers 100 layers # of Parameters
$m^{\prime}=32$ 92.95 $\pm$ 0.5 93.05 $\pm$ 0.3 93.64 $\pm$ 0.2 2,496
$m^{\prime}=64$ 91.95 $\pm$ 0.1 92.54 $\pm$ 0.1 92.85 $\pm$ 0.3 9,088
$m^{\prime}=128$ 91.20 $\pm$ 0.2 92.87 $\pm$ 0.1 93.24 $\pm$ 0.1 34,560
$m^{\prime}=256$ 90.73 $\pm$ 0.1 92.67 $\pm$ 0.2 92.90 $\pm$ 0.3 134,656
F.2 Extended experimental proofs on Over-smoothing and Over-Squashing mitigation
Here, we provide additional plots of Dirichlet Energy and layer-wise sensitivity to witness the model’s resilience to over-smoothing and over-squashing. Here we perform one training per model, sharing for each model the hyperparameters in Table 9. GCN is implemented such that over-smoothing is prone to occur, so we did not use residual connections and used ReLU as activation function [Scholkemper et al., 2024, Álvaro Arroyo et al., 2025].
@@ F.2 Extended experimental proofs on Over-smoothing and Over-Squashing mitigation
We report the sensitivity $S$ for Pascal-VOC and Peptides-struct in Figure 5. We notice that, similarly to the example in the main paper, we do not record any difficulties by the models to keep high sensitivity. Also, we notice that in Peptides-struct, we have a similar trend to Peptides-func, which reflects the same dataset domain. In Figure 5(a), we record a decreasing trend in sensitivity from all the models, but in general, they all have $S_{l}\neq 0$, also in this case. In Figure 6, we analyse the $E^{dir}$ trends for GCN, GRAFF, A-DGN, and SAS-GNN, to understand whether SAS-GNN effectively mitigates over-smoothing. We observe that also in this case, as in the Minesweeper example of Figure 2(a), SAS-GNN follows a similar trend of A-DGN, in all the other datasets, meaning that its ability to mitigate over-smoothing is more inherited from this model, rather than GRAFF. In these other experiments, we notice that $E^{dir}$ for GRAFF grows exponentially. As observed also in Minesweeper, it is clear that these plots describe the over-sharpening phenomenon discussed by Di Giovanni et al. [2023]. For instance, GRAFF is a more expressive version of GCN, which is provably capable of both asymptotically experiencing over-smoothing, as well as the opposite effect, namely over-sharpening. The desired behavior is a trade-off of the two, which is what we observe when GRAFF has a contained number of layers. Here, GRAFF underperforms due to an asymptotic behavior that lets adjacent node features diverge; indeed, $E^{dir}$ increases. As concerns GCN, as expected, it experiences over-smoothing also in the remaining heterophilic datasets.

### v2

@@ E.3 Implementation Details
The metrics that we use are the Area Under the Curve (AUC) for the binary classification experiments, accuracy for Roman Empire and Amazon Ratings, average precision (AP) for Peptides-func, mean-absolute error (MAE) for Peptides-struct, and macro F1-Score (F1) for Pascal-VOC. All the experiments concerning EEGNN and SAS-GNN implied that we removed self-loops from the graphs, in order to satisfy the assumptions made in Theorems 3.1, 3.2. We always apply $f(\mathbf{X})$ implemented via a one-layer perception with ReLU. And in the LRGB, we also use a final decoder MLP to refine the representations of $\mathbf{Z}$, details on the hyperparameter are in the code. In the heterophilic graphs, each training lasts 3000 epochs, and the optimizer that we use is Adam [Kingma and Ba, 2017]. Concerning the quantitative results, we report the mean and standard deviation, which are computed from the metric averaged across 10 runs. Each run corresponds to a different split of the nodes used for training, validation, and testing. These experiments were run on a single Nvidia GeForce RTX 3090 Ti 24 GB and also a single NVIDIA A100 80GB PCIe.
The experiments on ECHO, share same optimizers, metrics, and task objectives (except for APGCN which uses an additional loss term). There we run an hyperparameter optimization via optuna [Akiba et al., 2019] with 32 trials per model, and 1000 epochs per trial. The results we present are averaged across 4 weights initialization.
As concerns the LRGB experiments, we run each configuration for a maximum of 1000 epochs across 4 different weights initializations, using the AdamW optimizer. We picked the best configuration based on the validation metric. These experiments were run on a single NVIDIA A100 80GB PCIe.
As concerns with the experiments on the TUDatasets for short-range graph classification (see Table 25) and homophilic node classification (see Section F.9), we also used a single NVIDIA A100 80GB PCIe. All the data was taken mostly from the PyTorch Geometric library [Fey and Lenssen, 2019], and other open-source sources. We always refer to their provenance in the code when using them.
@@ E.4 Hyperparameters
For a fair comparison, the hyperparameter tuning process was done according to the hyperparameter set used by Finkelshtein et al. [2024]. In the SAS-GNN experiments, we choose $\tau$ among $\{0.1,0.5,1\}$. While in the experiments with EEGNNs, we do not need to tune $\tau$ because it is adaptive (see Algorithm 1), and we do not need to tune the number of layers parameter since we use a large and upper bound depth $L=20$. In the case of EEGNN, we choose $f_{c}(\cdot)$ as a Mean-GNN [Hamilton, 2020]. Similarly to [Finkelshtein et al., 2024], we test a contained number of layers $\{1,2,3\}$, and then as hidden dimension we select $4$, 88, $16$, $32$ or $64$. Then we also tune the value of the temperature $\nu_{0}$ between 0 and 0.1. $f_{\nu}(\cdot)$, the network that learns the adaptive temperature $\nu$, adopts the design proposed in [Finkelshtein et al., 2024]. In the ECHO benchmark, we performed the hyperparameter tuning based on the following hyperparameter ranges in Tables 9, 10, 11, 12. There we also consider the best values per-dataset.
Table 9: Hyperparameter search space and best configurations for EEGNN (Adastep) on ECHO tasks. $\tau$ and $L$ are fixed as they depends on node features at each layer.
Hyperparameter Search Space SSSP Eccentricy Diameter
$L$ 40 40 40 40
$m^{\prime}$ $[32,256]$ 69 57 254
$m^{\prime}_{f}$ $[32,128]$ 87 55 113
$L_{f}$ $[1,40]$ 1 14 24
Learning Rate $10^{-5},10^{-2}]$ 0.001907 0.006095 0.003307
Weight Decay $10^{-8},10^{-3}]$ 0.000919 0.000133 0.000744
$\tau$ 1.0 1.0 1.0 1.0
$L_{\nu}$ 2 2 2 2
$\nu_{0}$ 1.0 1.0 1.0 1.0
Table 10: Hyperparameter search space and best configurations for SAS-GNN on ECHO tasks.
Hyperparameter Search Space SSSP Eccentricity Diameter
$L$ $[1,40]$ 32 27 19
$m^{\prime}$ $[32,256]$ 50 196 66
Learning Rate $10^{-5},10^{-2}]$ 0.001907 0.000556 0.006141
Weight Decay $10^{-8},10^{-3}]$ 0.000270 0.000246 0.000165
$\tau$ $[0.01,1.0]$ 0.000165 0.09946355197568157 0.4037
Table 11: Hyperparameter search space and best configurations for GRAFF on ECHO tasks.
Hyperparameter Search Space SSSP Eccentricity Diameter
$L$ $[1,40]$ 21 15 9
$m^{\prime}$ $[32,256]$ 37 32 123
Learning Rate $10^{-5},10^{-2}]$ 0.007387 0.006607 0.003168
Weight Decay $10^{-8},10^{-3}]$ 0.000259 0.000134 0.000077
$\tau$ $[0.01,1.0]$ 0.3253 0.8447 0.8447
Self Loops False False False False
Table 12: Hyperparameter search space and best configurations for APGCN on ECHO tasks. Here $L$ is fixed, as we adapt the value at each layer, based on the node features. $L_{enc}$ stands for number of layers that the encoder of input node features has. $\mu$ is the multiplying coefficient of the budget-aware loss.
Hyperparameter Search Space SSSP Eccentricity
$L$ 40 40 40
$m^{\prime}$ $[32,256]$ 183 183
Learning Rate $10^{-5},10^{-2}]$ 0.002801 0.003853
Weight Decay $10^{-8},10^{-3}]$ 0.000672 0.000997
$L_{enc}$ 5 5 5
$\mu$ $10^{-4},10^{-2}]$ 0.0054 0.0069
In this case the exit and temperature networks, namely $f_{c}$ and $f_{\nu}$, were implemented using another SAS-GNN, so the overall parameter complexity can be simplified to $\mathcal{O}\!\left(m^{\prime 2}+m_{f}^{\prime 2}\right)$. In the experiments concerning the LRGB datasets, $f_{c}$ and $f_{\nu}$ keep the same hyperparameter set as the heterophilic experiments. The main hyperparameter set is the same used in [Tönshoff et al., 2023]. In any experiment with SAS-GNN and EEGNN, we do not use any dropout or normalization layer within the message-passing updates, to follow the ODE interpretation and directly minimize the energy functional in Equation (3). We only use them in the feature encoder and in the final decoder.
In the experiments with Peptides-func and Peptides-struct, we use two positional encoding techniques, which are typically approached in these settings. For Peptides-func we used RWSE [Dwivedi et al., 2022], while instead in Peptides-struct, we used LapPE [Dwivedi and Bresson, 2021b]. We used these according to the best hyperparameters used for graph convolutional methods in [Tönshoff et al., 2023]. We do not use positional encodings in Pascal-VOC.
@@ Appendix F Additional Results
In this section of the appendix we are going to include additional results that are incremental w.r.t. the claims made in the main paper. We show empirical validation of our theoretical properties, and also prove that both in synthetic and real-world experiments representative of long-range tasks and highly heterophilic graphs, SAS-GNN showcases its inductive bias, indicating that are good proxy to mitigate over-squashing and over-smoothing. We also provide complementary experiments that further support our claims or prove additional aspects, such as the scalability of our model.
We then move to provide a broader view on the real-world experiments presented in Tables 17 and 18. Then we also provide additional experiments on homophilic node classification, short-range graph classification on both medium-size and large scale datasets. Other content of this section are ablation and sensitivity studies on the early-exit components.
F.1 Empirical Validation of the SAS-GNN Theoretical Properties
Having introduced the new SAS-GNN backbone, and proved the theorem behind its architectural choices, we now turn to validating the theoretical claims underlying the SAS-GNN block. We already showed in the main text that SAS-GNN is capable of deep processing, and long-range reasoning. However, we want to understand whether this also translates in possible OST and OSQ mitigation, namely, do the SAS-GNN inductive biases also a good proxy to mitigate over-smoothing and over-squashing? We believe this might be the case according to the theoretical properties expressed in Section 3.
Lower values of $E^{dir}$ indicate smoothing; values near zero signal OST [Cai and Wang, 2020]. In Figure 7(a), we report $E^{dir}(\mathbf{H}^{t})$ at each layer on Minesweeper, a heterophilic dataset. GCN rapidly collapses, while GRAFF oversharpens, as described in Di Giovanni et al. [2023]. SAS-GNN and A-DGN, in contrast, maintain stable energy profiles, suggesting SAS-GNN inherits its stabilizing behavior primarily from A-DGN.
These trends affirm that SAS-GNN is effective as a proxy to contrast OST and OSQ. We must notice that the sensitivity metric do not directly distinguish among the types of OSQ, so we just show this as a way of highlighting the general advantage of SAS-GNN w.r.t. existing methods.
(a) $E^{dir}$ on Minesweeper.
[…]
We provide additional plots of Dirichlet Energy and layer-wise sensitivity to witness the model’s resilience to over-smoothing and over-squashing. Here we perform one training per model, sharing for each model the hyperparameters in Table 13. GCN is implemented such that OST is prone to occur, so we did not use residual connections and used ReLU as activation function [Scholkemper et al., 2024, Álvaro Arroyo et al., 2025].
@@ F.1 Empirical Validation of the SAS-GNN Theoretical Properties
We report the sensitivity $S$ for Pascal-VOC and Peptides-struct in Figure 8. We notice that, similarly to the example in the main paper, we do not record any difficulties by the models to keep high sensitivity. Also, we notice that in Peptides-struct, we have a similar trend to Peptides-func, which reflects the same dataset domain. In Figure 8(a), we record a decreasing trend in sensitivity from all the models, but in general, they all have $S_{l}\neq 0$, also in this case. In Figure 9, we analyse the $E^{dir}$ trends for GCN, GRAFF, A-DGN, and SAS-GNN, to understand whether SAS-GNN effectively mitigates over-smoothing. We observe that also in this case, as in the Minesweeper example of Figure 7(a), SAS-GNN follows a similar trend of A-DGN, in all the other datasets, meaning that its ability to mitigate over-smoothing is more inherited from this model, rather than GRAFF. In these other experiments, we notice that $E^{dir}$ for GRAFF grows exponentially. As observed also in Minesweeper, it is clear that these plots describe the over-sharpening phenomenon discussed by [Di Giovanni et al., 2023]. For instance, GRAFF is a more expressive version of GCN, which is provably capable of both asymptotically experiencing over-smoothing, as well as the opposite effect, namely over-sharpening. The desired behavior is a trade-off of the two, which is what we observe when GRAFF has a contained number of layers. Here, GRAFF underperforms due to an asymptotic behavior that lets adjacent node features diverge; indeed, $E^{dir}$ increases. As concerns GCN, as expected, it experiences over-smoothing also in the remaining heterophilic datasets.
@@
F.2 SAS-GNN is able to retain performance as depth increases
Having proposed SAS-GNN, we have demonstrated that it is always stable and non-dissipative, and also able to allow each node to induce attraction or repulsion edge-wise. We study the behavior of SAS-GNN when the depth increases up to 50 layers, which is an overestimate of what we need in practice. In our experiments, we design SAS-GNN, as specified in Corollary 3.3. All the data points corresponds to a different training procedure, where we display the best accuracy. We present the results in Figure 10(a), where we show an experiment with Cora, a homophilic dataset. In Figure 10(b), we assess the resilience to depth via Minesweeper, an example of a heterophilic graph. In Cora, we notice that both A-DGN and SAS-GNN can retain their performance. Also, GRAFF can maintain it up to the 30th layer, then it starts decreasing. This behavior is expected since GRAFF is only designed to contrast over-smoothing, which does not occur, but it is not specifically designed to be robust to the number of layers. It is interesting to notice that SAS-GNN, even though part of its architecture is inherited by GRAFF, can retain its performance, behaving similarly to A-DGN. As far as GCN is concerned, we see that it has started losing its performance in the earlier layers. Here, we consider a GCN design that lacks the residual connection and does not use weight sharing. As concerns Minesweeper, we notice that GRAFF is the model that maximizes the most $E^{dir}$, while SAS-GNN and A-DGN still have a similar trend. None of these models seems to experience over-smoothing, except for GCN, where $E^{dir}$ approaches 0. Also, in this case, all the models retain their performances, except for GCN, which starts decreasing from the beginning. Since GRAFF is the only model that perceive the performance degradation, and Dirichlet Energy does not record over-smoothing, we conclude that it probably suffer from other phenomena, that do not affect A-DGN and SAS-GNN.
Table 14: Performance across different model sizes and layer depths
Model 10 layers 50 layers 100 layers # of Parameters
$m^{\prime}=32$ 92.95 $\pm$ 0.5 93.05 $\pm$ 0.3 93.64 $\pm$ 0.2 2,496
$m^{\prime}=64$ 91.95 $\pm$ 0.1 92.54 $\pm$ 0.1 92.85 $\pm$ 0.3 9,088
$m^{\prime}=128$ 91.20 $\pm$ 0.2 92.87 $\pm$ 0.1 93.24 $\pm$ 0.1 34,560
$m^{\prime}=256$ 90.73 $\pm$ 0.1 92.67 $\pm$ 0.2 92.90 $\pm$ 0.3 134,656
We also provide further experiments showing that SAS-GNN preserves its predictive power even when the number of layers increases. This property is crucial for two reasons. First, deeper models enable richer feature extraction, which can improve generalization, something that standard MPNNs often fail to achieve due to their limitations. Second, in the specific case of MPNNs, increased depth allows for long-range information propagation, which shallower networks cannot capture, but could be critical in some settings.
To assess these two aspects, we examine two different scenarios:
1. Highly heterophilic graphs. In this setting, the challenge lies not in long-range propagation but in the ability to correctly separate adjacent nodes that frequently belong to different classes. We focus on the WebKB datasets introduced by García-Plaza et al. [2016], namely Texas, Wisconsin, and Cornell. Their statistics are reported in Table 15.
Table 15: Highly-Heterophilic Graphs.
Datasets NN $|\mathcal{E}|$ $d$ $|C|$ $\xi_{edge}$ $\xi_{adj}$
Texas 183 574 1703 5 0.09 -0.26
Wisconsin 251 916 1703 5 0.20 -0.15
Cornell 183 557 1703 5 0.13 -0.21
Although these datasets are no longer widely used for benchmarking [Platonov et al., 2023], since models typically exhibit high variance across folds, they are valuable examples of extreme heterophily. Both $\xi_{edge}$ and $\xi_{adj}$ take very low values, meaning that neighbors share different classes very frequently, even regardless of class distribution (i.e. $\xi_{edge}$ does not consider number of classes of class balance). Compared to the benchmarks introduced by Platonov et al. [2023], the WebKB graphs display significantly stronger heterophily.
For this scenario, depth is not used to capture long-range dependencies but rather as a stress test: we want to see whether the model can maintain performance across layers while still separating adjacent nodes correctly. SAS-GNN is a hybrid of GRAFF, which addresses over-smoothing and heterophily, and A-DGN, which addresses over-squashing and long-range propagation. Thus, it is particularly interesting to evaluate its behavior here. In the main paper, we already showed that SAS-GNN is actually effective in this scenario by maintaining its performance on Texas even as depth increases (see Figure 1), whereas classic MPNNs, even with residual connections, which safeguard from over-smoothing [Scholkemper et al., 2024], still suffer performance drops at larger depths. In Figure 11, we report additional results on Cornell and Wisconsin, confirming the same trend. The common aspects among these comparison are that GRAFF is not able to retain its performance across layers, which is expected, since it was not designed for deep configurations. Specifically, in the original paper of GRAFF, the best configurations were typically using very shallow configurations (i.e. 2-3 layers [Di Giovanni et al., 2023]). The other aspect is that SAS-GNN is always retaining accuracy, showing also that regardless of the hostility of the class distribution, it still preserve information, similarly to A-DGN. A-DGN seems to lag behind SAS-GNN in these scenarios, probably because A-DGN was not designed to deal with high heterophily. Nonetheless, we did not tune extensively on of these models, these experiments are mostly a proof of concept that illustrates the capabilities of our model and how its inductive biases reflect on real-world graphs. The details to reproduce these experiments are in the codebase.
(c) Wisconsin
(d) Cornell
Figure 11: Accuracy comparison of all baselines on Wisconsin and Cornell.
2. Synthetic datasets for long-range propagation. In contrast, our second scenario is designed to explicitly test whether depth enables the model to exploit long-range information. For this purpose, we use synthetic graphs where the predictive task depends on aggregating information across distant nodes. In this setting, the ability of SAS-GNN to propagate signals over many layers is the main factor under evaluation, rather than heterophily and retaining performance after many layers.
Following the experimental setup of Gravina et al. [2025], we consider the Graph Transfer task. This resilience of a model to over-squashing when transmitting information across multiple hops.
Graph Transfer.
The Graph Transfer task focuses directly on the problem of topological and computational over-squashing. The setup consists of a source–target pair of nodes separated by $k$ hops ($k\in\{3,5,10,50\}$). The goal is to transfer a binary label from the source to the target across different graph topologies (line, ring, and crossed-ring). A better definition of the task is described by Gravina et al. [2025]. This task can only be solved by preserving source information over long distances, where topology may be responsible for compressing excessively such information.
In Figure 12, we report the results of our experiments. SAS-GNN maintains stable performance across depths, confirming that it inherits the inductive bias of A-DGN, which was explicitly designed for long-range propagation. However, SAS-GNN performs slightly below A-DGN and SWAN in this benchmark. Notably, GRAFF alone performs poorly in this setting, and this suggests that the symmetric component inherited from it does not provide benefits here and likely reduces predictive accuracy. This likely explains why SAS-GNN lags behind the other two models despite retaining stability across layers.
Figure 12: Performance on the Graph Transfer task for line, ring, and crossed-ring graphs at increasing distances. SAS-GNN remains stable across depths, though it lags behind A-DGN.
Summary.
Taken together, these results show that SAS-GNN is effective under multiple challenging conditions: extreme heterophily, very deep architectures, and long-range propagation. By combining the strengths of GRAFF (robustness to heterophily) and A-DGN (resilience to over-squashing and long-range propagation), SAS-GNN emerges as a reliable architecture for both deep processing and information preservation. This makes it a strong candidate not only for tackling over-smoothing and over-squashing simultaneously, but also for practical applications such as the design of early-exit strategies.